# Ringdown in Vaidya spacetimes: time-dependent frequencies, Penrose limit and time-domain analyses

### v2

@@ I. Introduction
The detections of the gravitational waves from a binary black hole [1, 2] have opened a new window onto the black holes in our universe. The signal from compact object binaries typically consists of three phases: inspiral, merger, and ringdown. The ringdown phase, characterized by a superposition of damped sinusoids known as quasinormal modes (QNMs) [3, 4], is particularly significant as it represents the final perturbed black hole settling down to a Kerr black hole (see, e.g., reviews [5, 6, 7]). The frequencies and damping rates of these QNMs are uniquely determined by the mass and spin of the final Kerr black hole, offering a powerful tool for “black hole spectroscopy”, which tests the predictions of General Relativity measuring the QNMs of the remnant  [8, 9, 10].
@@
Intriguingly, a deep connection has been established between the properties of QNMs in the eikonal (high frequency, geometric optics) limit and the dynamics of null geodesics near the black hole photon sphere (or photon ring) in static or stationary spacetimes [11, 12, 13]. Specifically, the real part of the QNM frequency ($\omega_{\rm Re}$) is related to the angular velocity ($\Omega$) of unstable circular null orbits, while the imaginary part ($\omega_{\rm Im}$), representing the damping rate, is linked to the Lyapunov exponent ($\lambda$) quantifying the instability of these orbits along the radial direction. The correspondence typically takes the form $\omega\approx\Omega\ell-i\lambda(n+1/2)$ for large angular momentum quantum number $\ell$ and overtone number $n$. This connection provides valuable physical intuition for the origin of ringdown radiation, associating it with photons (or gravitons) temporarily trapped in unstable orbits near the black hole before escaping to infinity or falling into the black hole. This correspondence—referred to as the QNM-geodesic correspondence–has been checked in various cases of static, vacuum black holes in 4 and higher dimensions [13]. (See e.g., [14] for the non-vacuum cases, and e.g., [15] for further studies). It is of considerable interest to study whether or not such a correspondence holds in more generic, dynamical cases.
@@
The Penrose limit [16] offers a powerful mathematical framework to explore this connection further. By focusing on the geometry in the infinitesimal vicinity of a null geodesic, the Penrose limit transforms the spacetime metric into a simpler plane-wave (or pp-wave) form [17]. When applied to an unstable circular orbit of photons (UCOP) in a static, spherically symmetric spacetime, the resulting Penrose limit metric inherits the key dynamical properties of the orbit. Analyzing wave propagation (e.g., a scalar field) within this simplified geometry naturally reproduces the eikonal QNM spectrum, with the metric components directly related to the orbital frequency $\Omega$ and the Lyapunov exponent $\lambda$, thereby solidifying the QNM-geodesic correspondence [18, 13, 19, 20]. Recent developments of QNM analyses can be seen in, e.g., Refs. [21, 22, 23, 24, 25, 26].
@@
However, most astrophysical black holes are not isolated or perfectly static; they interact with their environment, accreting matter or potentially radiating energy. This raises the fundamental question of how QNMs and the ringdown process can be characterized in dynamical spacetimes. The Vaidya spacetime [27], describing a spherically symmetric radiating or accreting object composed of null dust, provides a simple yet non-trivial model for studying such situations (see Refs. [28, 29, 30, 31, 32, 33, 34] for previous works on ringdown waves in dynamical spacetimes). Recent work has shown that the Vaidya spacetime possesses a dynamical photon sphere, whose radius evolves in response to the changing mass function [35].
@@
In this paper, we extend the investigation of the QNM-geodesic correspondence to the dynamical context of the Vaidya spacetime. Our primary tool is the Penrose limit, adapted not to a static UCOP, but to a null geodesic lying on the dynamical photon sphere trajectory identified in [35]. By calculating the Penrose limit geometry along this evolving trajectory, we extract time-dependent counterparts of the orbital frequency and Lyapunov exponent within an adiabatic approximation. We hypothesize that these quantities characterize the temporal properties of ringdown waves emitted during the dynamical phase. To test this hypothesis, we perform numerical simulations of tensor perturbations propagating in the full Vaidya spacetime during both constant and time-dependent accretion phases. We then extract the time-varying frequency and damping rate from the resulting ringdown waveform and compare these numerical results with the predictions derived from our Penrose limit analysis. This comparison allows us to assess the extent to which the Penrose limit geometry around the dynamical photon sphere captures the essential features of ringdown in this non-stationary background.
@@
This paper is organized as follows. In Sec. 2, we review the connection between QNMs and the Penrose limit for a static, spherically symmetric spacetime. In Sec. 3, we analyze the case of a Vaidya spacetime with a constant accretion rate, comparing results from the Penrose limit, frequency-domain QNM calculations, and time-domain simulations. In Sec. 4, we extend this analysis to Vaidya spacetimes with time-dependent accretion rates, comparing the predictions from the adiabatic Penrose limit with full numerical ringdown simulations. Section 5 provides a summary and discussion of our findings. We use geometric units where $G=c=1$ throughout the paper.
@@ A. Unstable circular orbits of photons
Since we are interested in the Penrose limit geometry adapted to an unstable circular orbit of photons (UCOP), let us review the null geodesics and UCOP in a static spherically symmetric spacetime in this subsection. Let us write the metric of a static spherically symmetric spacetime as follows:
$\mathrm{d}s^{2}=-f(r)\mathrm{d}t^{2}+h(r)\mathrm{d}r^{2}+r^{2}\mathrm{d}\theta^{2}+r^{2}\sin^{2}\theta\mathrm{d}\phi^{2}.$ (1)
Because of the symmetry, we can focus on the equatorial surface, and in what follows we set $\theta=\pi/2$ without loss of generality. Then the Lagrangian for null geodesics is given by
$\mathcal{L}=\frac{1}{2}\left(-f\dot{t}^{2}+h\dot{r}^{2}+r^{2}\dot{\phi}^{2}\right),$ (2)
@@
$\dot{t}$ $=$ $\frac{E}{f},$ (3)
$\dot{\phi}$ $=$ $\frac{L}{r^{2}},$ (4)
$0$ $=$ $\frac{1}{2}\dot{r}^{2}+V(r),$ (5)
@@
$V^{\prime\prime}(r_{\rm c})=\frac{L^{2}}{2h_{\rm c}r_{\rm c}^{2}}\left(-\frac{2}{r_{\rm c}^{2}}+\frac{f^{\prime\prime}_{\rm c}}{f_{\rm c}}\right),$ (12)
@@
This circular orbit is unstable against small deviations along the radial direction. The unstable behavior can be characterized by the characteristic exponent $\lambda$ as $\sim{\rm e}^{\lambda t}$. Substituting the functional form of the small deviation $\delta r\sim{\rm e}^{\lambda t}$ into Eq. (8), we obtain
$\lambda^{2}=-\frac{V^{\prime\prime}_{\rm c}}{\dot{t}|_{r=r_{\rm c}}^{2}}=-V^{\prime\prime}_{\rm c}\frac{f_{\rm c}^{2}}{E^{2}}.$ (13)
@@
$\Omega:=\left.\frac{\mathrm{d}}{\mathrm{d}t}\phi\right|_{r=r_{\rm c}}=\left.\frac{\dot{\phi}}{\dot{t}}\right|_{r=r_{\rm c}}=\frac{L/r_{\rm c}^{2}}{E/f_{\rm c}}=\frac{\sqrt{f_{\rm c}}}{r_{\rm c}}.$ (14)
@@ B. Coordinate transformation and Penrose limit
Here we derive the Penrose limit metric in an ad hoc way, for the sake of a simple argument. The general derivation of the Penrose limit is shown in Appendix A.1, and the derivation of the Penrose coordinates adapted to a UCOP is described in Appendix B. First, let us express a UCOP in a parametric form by using the affine parameter $\tau$ as
$t$ $=$ $\frac{E}{f_{\rm c}}\tau,$ (15)
$r$ $=$ $r_{\rm c},$ (16)
$\theta$ $=$ $\frac{\pi}{2},$ (17)
$\phi$ $=$ $\frac{L}{r_{\rm c}^{2}}\tau=\frac{E}{r_{\rm c}\sqrt{f_{\rm c}}}\tau=:\beta\tau.$ (18)
@@
$t$ $=$ $\frac{E}{f_{\rm c}}(u+\varepsilon^{2}\Lambda v)-\frac{\varepsilon^{2}}{E}v,$ (19)
$r$ $=$ $r_{\rm c}+\frac{\varepsilon}{\sqrt{h_{\rm c}}}x,$ (20)
$\theta$ $=$ $\frac{\pi}{2}+\frac{\varepsilon}{r_{\rm c}}y,$ (21)
$\phi$ $=$ $\beta(u+\varepsilon^{2}\Lambda v),$ (22)
@@
$\mathrm{d}\bar{s}^{2}$ $=$ $\lim_{\varepsilon\rightarrow 0}\left(\varepsilon^{-2}\left.\mathrm{d}s^{2}\right|_{(t,r,\theta,\phi)\rightarrow(u,v,x,y)}\right),$ (23)
$=$ $\mathrm{d}u\mathrm{d}v+(\alpha^{2}x^{2}-\beta^{2}y^{2})\mathrm{d}u^{2}+\mathrm{d}x^{2}+\mathrm{d}y^{2},$ (24)
where $\alpha=-V^{\prime\prime}_{\rm c}$. This is the Penrose limit metric adapted to a UCOP in the Brinkmann coordinates. The metric components $\alpha$ and $\beta$ are related to $\Omega$ and $\lambda$ as
$\alpha$ $=$ $\dot{t}|_{r=r_{\rm c}}\lambda=\frac{E}{f_{\rm c}}\lambda,$ (25)
$\beta$ $=$ $\dot{t}|_{r=r_{\rm c}}\Omega=\frac{E}{f_{\rm c}}\Omega.$ (26)
@@
C. Quasinormal modes in the Penrose limit
Let us consider a massless scalar field $\Psi$ in the Penrose limit metric (24). In this section, we basically follow the discussion in Ref. [20]. The wave equation is given by
$\nabla_{\mu}\nabla^{\mu}\Psi=\left[2\partial_{u}\partial_{v}-\left(\alpha^{2}x^{2}-\beta^{2}y^{2}\right)\partial_{v}^{2}+\partial_{x}^{2}+\partial_{y}^{2}\right]\Psi=0.$ (27)
@@ C. Quasinormal modes in the Penrose limit
$\Psi={\rm e}^{-\mathrm{i}\tilde{\omega}u+\mathrm{i}pv}X(x)Y(y),$ (28)
@@
$\mu_{x}X$ $=$ $-\left(\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+p^{2}\alpha^{2}x^{2}\right)X,$ (29)
$\mu_{y}Y$ $=$ $-\left(\frac{\mathrm{d}^{2}}{\mathrm{d}y^{2}}-p^{2}\beta^{2}y^{2}\right)Y,$ (30)
$0$ $=$ $2\tilde{\omega}p-\mu_{x}-\mu_{y}.$ (31)
@@
$\mu_{y}=2\beta p\left(n_{y}+\frac{1}{2}\right),$ (32)
where $n_{y}$ takes a non-negative integer. For the IHO, we are interested in the pure outgoing solutions, which can be obtained just replacing the parameter $p$ by $\mathrm{i}p$, namely,
$\mu_{x}=2\mathrm{i}\alpha p\left(n_{x}+\frac{1}{2}\right)$ (33)
@@
$\tilde{\omega}=\beta\left(n_{y}+\frac{1}{2}\right)+\mathrm{i}\alpha\left(n_{x}+\frac{1}{2}\right).$ (34)
In the standard way of the separation of variables with respect to the original coordinates $(t,r,\theta,\phi)$, the wave function is proportional to ${\rm e}^{-\mathrm{i}\omega t+\mathrm{i}m\phi}$ with the frequency $\omega$ and a magnetic quantum number $m$. Rewriting this factor in terms of the coordinates $u$ and $v$ through Eqs. (19) and (22), we obtain
$\exp\left[-\mathrm{i}\omega t+\mathrm{i}m\phi\right]=\exp\left[\mathrm{i}\left(\frac{\omega}{E}-\frac{E\omega}{f_{\rm c}}\Lambda+m\beta\Lambda\right)v-\mathrm{i}\left(\frac{L^{2}}{Er_{\rm c}^{2}}\omega-\beta m\right)u\right],$ (35)
where we have reset the bookkeeping parameter $\varepsilon$ to $1$, or in other words, renormalized the coordinates $v$ including $\varepsilon^{2}$. Hereafter, we also renormalize the coordinates $x$ and $y$ by simply setting $\varepsilon=1$. Since the exponent $\mathrm{i}\omega t+\mathrm{i}m\phi$ must be equal to $\mathrm{i}\tilde{\omega}u+\mathrm{i}pv$ in the Penrose limit, we obtain the following relations:
$p$ $=$ $\frac{\omega}{E}-\frac{E\omega}{f_{\rm c}}\Lambda+m\beta\Lambda,$ (36)
$\tilde{\omega}$ $=$ $\beta\left(\frac{L}{E}\omega-m\right).$ (37)
@@
$\omega$ $=$ $\frac{E}{L}\left(n_{y}+m+\frac{1}{2}\right)+\mathrm{i}\frac{E}{L}\frac{\alpha}{\beta}\left(n_{x}+\frac{1}{2}\right)$ (38)
$=$ $\Omega\left(n_{y}+m+\frac{1}{2}\right)+\mathrm{i}\lambda\left(n_{x}+\frac{1}{2}\right).$ (39)
Therefore, identifying $n_{y}+m$ to $\ell$ we can establish the correspondence between the null geodesics and QNMs reported in Ref. [13]. Since $\omega$ is a complex variable, $p$ is also complex in general, and the amplitude of $\Psi$ changes along the $v$-direction unless $p$ is a real value. To make the value of $p$ real, we need to fix the arbitrary constant $\Lambda$ to be $\frac{f_{\rm c}}{E^{2}}$. Then we obtain $p=m/L$ and the relevant components of the coordinate transformation become
$t$ $=$ $\frac{E}{f_{\rm c}}u,$ (40)
$\phi$ $=$ $\beta u+\frac{v}{L}.$ (41)
@@
$P^{\ell}_{\ell}(\cos\theta)\sim\left(1-\cos^{2}\left(\frac{\pi}{2}+\frac{y}{r_{\rm c}}\right)\right)^{\ell/2},$ (42)
@@
$P^{\ell}_{\ell}(\cos\theta)\sim 1-\frac{\ell}{2}\frac{y^{2}}{r_{\rm c}^{2}}.$ (43)
A similar expression can be obtained by expanding the ground state for the SHO described by Eq. (30) as follows:
$\exp\left[-\frac{1}{2}\beta py^{2}\right]\sim\exp\left[-\frac{1}{2}\beta\frac{m}{L}y^{2}\right]=\exp\left[-\frac{1}{2}\frac{mE}{f_{\rm c}L}\Omega y^{2}\right]\sim 1-\frac{n_{y}+m}{2}\frac{y^{2}}{r_{\rm c}^{2}},$ (44)
@@
$\partial_{\theta}+\mathrm{i}\cot\theta\partial_{\phi}\sim-r_{\rm c}\partial_{y}-\cot\left(\frac{y}{r_{\rm c}}+\frac{\pi}{2}\right)m\sim-r_{\rm c}\partial_{y}+\frac{L}{r_{\rm c}}py=r_{\rm c}(-\partial_{y}+\beta py).$ (45)

### v3

@@ I. Introduction
The detections of the gravitational waves from a binary black hole [1, 2] have opened a new window onto the black holes in our universe. The signal from compact object binaries typically consists of three phases: inspiral, merger, and ringdown. The ringdown phase, characterized by a superposition of damped sinusoids known as quasinormal modes (QNMs) [3, 4], is particularly significant as it represents the final perturbed black hole settling down to a Kerr black hole (see, e.g., reviews [5, 6, 7]). The frequencies and damping rates of these QNMs are uniquely determined by the mass and spin of the final Kerr black hole under the assumption that General Relativity is correct, the effects of surrounding matter fields are negligible, and hence the Kerr hypothesis holds. Therefore, measurement of the QNM frequencies offers a powerful tool for “black hole spectroscopy”, which tests the predictions of General Relativity   [8, 9, 10].
@@
Intriguingly, a deep connection has been established between the properties of QNMs in the eikonal (high frequency, geometric optics) limit and the dynamics of null geodesics near the black hole photon sphere (or photon ring) in static or stationary spacetimes [11, 12, 13]. Specifically, the real part of the QNM frequency ($\omega_{\rm Re}$) is related to the angular velocity ($\Omega$) of unstable circular null orbits, while the imaginary part ($\omega_{\rm Im}$), representing the damping rate, is linked to the Lyapunov exponent ($\lambda$) quantifying the instability of these orbits along the radial direction. The correspondence typically takes the form $\omega\approx\Omega\ell-\mathrm{i}\lambda(n+1/2)$ for large angular momentum quantum number $\ell$ and overtone number $n$. This connection provides valuable physical intuition for the origin of ringdown radiation, associating it with photons (or gravitons) temporarily trapped in unstable orbits near the black hole before escaping to infinity or falling into the black hole. This correspondence—referred to as the QNM-geodesic correspondence–has been checked in various cases of stationary, vacuum black holes in 4 and higher dimensions [13]. (See e.g., [14] for the non-vacuum cases, and e.g., [15] for further studies.) It is of considerable interest to study whether or not such a correspondence holds in more generic, dynamical cases.
@@
The Penrose limit (PL) [16] offers a powerful mathematical framework to explore this connection further. By focusing on the geometry in the infinitesimal vicinity of a null geodesic, the PL transforms the spacetime metric into a simpler plane-wave (or pp-wave) form [17]. When applied to an unstable circular orbit of photons (UCOP) in a stationary, spherically symmetric spacetime, the resulting PL metric inherits the key dynamical properties of the orbit. Analyzing wave propagation (e.g., a scalar field) within this simplified geometry naturally reproduces the eikonal QNM spectrum, with the metric components directly related to the orbital frequency $\Omega$ and the Lyapunov exponent $\lambda$, thereby solidifying the QNM-geodesic correspondence [18, 13, 19, 20]. Recent developments of QNM analyses can be seen in, e.g., Refs. [21, 22, 23, 24, 25, 26].
@@
However, most astrophysical black holes are not isolated or perfectly stationary; they interact with their environment, accreting matter or potentially radiating energy. For instance, a black hole formed soon after the neutron star merger would be surrounded by baryonic matter [2]. Furthermore, accretion of dark matter onto a black hole may not be ruled out with electromagnetic observations. From a quantum perspective, the black hole will continue to emit Hawking radiation, although this may not be relevant from an astrophysical perspective (see e.g., Ref. [27] for calculation of ringdown waveform in an evaporating black hole). This raises the fundamental question of how QNMs and the ringdown process can be characterized in dynamical spacetimes. The Vaidya spacetime [28], describing a spherically symmetric radiating or accreting object composed of null dust, provides a simple yet non-trivial model for studying such situations (see Refs. [27, 29, 30, 31, 32, 33, 34] for previous works on ringdown waves in dynamical spacetimes). Recent work has shown that the Vaidya spacetime possesses a dynamical photon sphere, whose radius evolves in response to the changing mass function [35].
@@
In this paper, we extend the investigation of the QNM-geodesic correspondence to the dynamical context of the Vaidya spacetime. Our primary tool is the PL adapted to a null geodesic lying on the dynamical photon sphere trajectory identified in [35]. By calculating the PL geometry along this evolving trajectory, we extract time-dependent counterparts of the orbital frequency and Lyapunov exponent within an adiabatic approximation. We hypothesize that these quantities characterize the temporal properties of ringdown waves emitted during the dynamical phase. To test this hypothesis, we perform numerical simulations of tensor perturbations propagating in the full Vaidya spacetime during both constant and time-dependent accretion phases. We then extract the time-varying frequency and damping rate from the resulting ringdown waveform and compare these numerical results with the predictions derived from our PL analysis. This comparison allows us to assess the extent to which the PL geometry around the dynamical photon sphere captures the essential features of ringdown in this non-stationary background.
@@
This paper is organized as follows. In Sec. 2, we review the connection between QNMs and the PL for a static, spherically symmetric spacetime. In Sec. 3, we analyze the case of a Vaidya spacetime with a constant accretion rate, comparing results from the PL, frequency-domain QNM calculations, and time-domain simulations. In Sec. 4, we extend this analysis to Vaidya spacetimes with time-dependent accretion rates, comparing the predictions from the adiabatic PL with full numerical ringdown simulations. Section 5 provides a summary and discussion of our findings. We use geometric units where $G=c=1$ throughout the paper.
@@ A. Unstable circular orbits of photons
Since we are interested in the PL geometry adapted to a UCOP, let us review the null geodesics and UCOP in a static spherically symmetric spacetime in this subsection. Let us write the metric of a static spherically symmetric spacetime as follows:
$\mathrm{d}s^{2}=-f(r)\mathrm{d}t^{2}+h(r)\mathrm{d}r^{2}+r^{2}\mathrm{d}\theta^{2}+r^{2}\sin^{2}\theta\mathrm{d}\phi^{2}.$ (2.1)
Because of the symmetry, we can focus on the equatorial surface, and in what follows we set $\theta=\pi/2$ without loss of generality. Then the Lagrangian for a free particle motion is given by
$\mathcal{L}=\frac{1}{2}\left(-f\dot{t}^{2}+h\dot{r}^{2}+r^{2}\dot{\phi}^{2}\right),$ (2.2)
@@
$\dot{t}$ $=$ $\frac{E}{f},$ (2.3)
$\dot{\phi}$ $=$ $\frac{L}{r^{2}},$ (2.4)
$0$ $=$ $\frac{1}{2}\dot{r}^{2}+V(r),$ (2.5)
@@
$V^{\prime\prime}(r_{\rm c})=\frac{L^{2}}{2h_{\rm c}r_{\rm c}^{2}}\left(-\frac{2}{r_{\rm c}^{2}}+\frac{f^{\prime\prime}_{\rm c}}{f_{\rm c}}\right),$ (2.12)
@@
This circular orbit is unstable against small deviations along the radial direction. The unstable behavior can be characterized by the characteristic exponent $\lambda$ as $\sim{\rm e}^{\lambda t}$. Substituting the functional form of the small deviation $\delta r\sim{\rm e}^{\lambda t}$ into Eq. (2.8), we obtain
$\lambda^{2}=-\frac{V^{\prime\prime}_{\rm c}}{\dot{t}|_{r=r_{\rm c}}^{2}}=-V^{\prime\prime}_{\rm c}\frac{f_{\rm c}^{2}}{E^{2}}.$ (2.13)
@@
$\Omega:=\left.\frac{\mathrm{d}}{\mathrm{d}t}\phi\right|_{r=r_{\rm c}}=\left.\frac{\dot{\phi}}{\dot{t}}\right|_{r=r_{\rm c}}=\frac{L/r_{\rm c}^{2}}{E/f_{\rm c}}=\frac{\sqrt{f_{\rm c}}}{r_{\rm c}}.$ (2.14)
@@ B. Coordinate transformation and Penrose limit
Here we derive the PL metric in an ad hoc way, for the sake of a simple argument. The general derivation of the PL is shown in Appendix A.1, and the derivation of the Penrose coordinates adapted to a UCOP is described in Appendix B. First, let us express a UCOP in a parametric form by using the affine parameter $\tau$ as
$t$ $=$ $\frac{E}{f_{\rm c}}\tau,$ (2.15)
$r$ $=$ $r_{\rm c},$ (2.16)
$\theta$ $=$ $\frac{\pi}{2},$ (2.17)
$\phi$ $=$ $\frac{L}{r_{\rm c}^{2}}\tau=\frac{E}{r_{\rm c}\sqrt{f_{\rm c}}}\tau=:\beta\tau.$ (2.18)
@@
$t$ $=$ $\frac{E}{f_{\rm c}}(u+\varepsilon^{2}\Lambda v)-\frac{\varepsilon^{2}}{E}v,$ (2.19)
$r$ $=$ $r_{\rm c}+\frac{\varepsilon}{\sqrt{h_{\rm c}}}x,$ (2.20)
$\theta$ $=$ $\frac{\pi}{2}+\frac{\varepsilon}{r_{\rm c}}y,$ (2.21)
$\phi$ $=$ $\beta(u+\varepsilon^{2}\Lambda v),$ (2.22)
@@
$\mathrm{d}\bar{s}^{2}$ $=$ $\lim_{\varepsilon\rightarrow 0}\left(\varepsilon^{-2}\left.\mathrm{d}s^{2}\right|_{(t,r,\theta,\phi)\rightarrow(u,v,x,y)}\right),$ (2.23)
$=$ $\mathrm{d}u\mathrm{d}v+(\alpha^{2}x^{2}-\beta^{2}y^{2})\mathrm{d}u^{2}+\mathrm{d}x^{2}+\mathrm{d}y^{2},$ (2.24)
where $\alpha=-V^{\prime\prime}_{\rm c}$. This is the PL metric adapted to a UCOP in the Brinkmann coordinates. The coefficients $\alpha$ and $\beta$ are related to $\Omega$ and $\lambda$ as
$\alpha$ $=$ $\dot{t}|_{r=r_{\rm c}}\lambda=\frac{E}{f_{\rm c}}\lambda,$ (2.25)
$\beta$ $=$ $\dot{t}|_{r=r_{\rm c}}\Omega=\frac{E}{f_{\rm c}}\Omega.$ (2.26)
@@
C. Quasinormal modes in the PL
Let us consider a massless scalar field $\Psi$ in the PL metric (2.24). In this section, we basically follow the discussion in Ref. [20]. The wave equation is given by
$\nabla_{\mu}\nabla^{\mu}\Psi=\left[2\partial_{u}\partial_{v}-\left(\alpha^{2}x^{2}-\beta^{2}y^{2}\right)\partial_{v}^{2}+\partial_{x}^{2}+\partial_{y}^{2}\right]\Psi=0.$ (2.27)
@@ C. Quasinormal modes in the PL
$\Psi={\rm e}^{-\mathrm{i}\tilde{\omega}u+\mathrm{i}pv}X(x)Y(y),$ (2.28)
@@
$\mu_{x}X$ $=$ $-\left(\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+p^{2}\alpha^{2}x^{2}\right)X,$ (2.29)
$\mu_{y}Y$ $=$ $-\left(\frac{\mathrm{d}^{2}}{\mathrm{d}y^{2}}-p^{2}\beta^{2}y^{2}\right)Y,$ (2.30)
$0$ $=$ $2\tilde{\omega}p-\mu_{x}-\mu_{y}.$ (2.31)
@@
$\mu_{y}=2\beta p\left(n_{y}+\frac{1}{2}\right),$ (2.32)
where $n_{y}$ takes a non-negative integer. For the IHO, we are interested in the pure outgoing solutions, which can be obtained just replacing the parameter $p$ by $\mathrm{i}p$, [^2] The IHO equation can be obtained by $p\rightarrow\pm\mathrm{i}p$. Then, including the time dependence, the asymptotic form of the eigen function transforms as $\exp\left[-\mathrm{i}\tilde{\omega}u-\alpha px^{2}/2\right]\rightarrow\exp\left[-\mathrm{i}(\tilde{\omega}u\pm\alpha px^{2}/2)\right]$. Therefore the lower sign corresponds to the pure outgoing boundary condition. namely,
$\mu_{x}=-2\mathrm{i}\alpha p\left(n_{x}+\frac{1}{2}\right)$ (2.33)
@@
$\tilde{\omega}=\beta\left(n_{y}+\frac{1}{2}\right)-\mathrm{i}\alpha\left(n_{x}+\frac{1}{2}\right).$ (2.34)
In the standard way of the separation of variables with respect to the original coordinates $(t,r,\theta,\phi)$, the wave function is proportional to ${\rm e}^{-\mathrm{i}\omega t+\mathrm{i}m\phi}$ with the frequency $\omega$ and a magnetic quantum number $m$. Rewriting this factor in terms of the coordinates $u$ and $v$ through Eqs. (2.19) and (2.22), we obtain
$\exp\left[-\mathrm{i}\omega t+\mathrm{i}m\phi\right]=\exp\left[\mathrm{i}\left(\frac{\omega}{E}-\frac{E\omega}{f_{\rm c}}\Lambda+m\beta\Lambda\right)v-\mathrm{i}\left(\frac{L^{2}}{Er_{\rm c}^{2}}\omega-\beta m\right)u\right],$ (2.35)
where we have reset the bookkeeping parameter $\varepsilon$ to $1$, or in other words, renormalized the coordinates $v$ including $\varepsilon^{2}$. Hereafter, we also renormalize the coordinates $x$ and $y$ by simply setting $\varepsilon=1$. Since the exponent $\mathrm{i}\omega t+\mathrm{i}m\phi$ must be equal to $\mathrm{i}\tilde{\omega}u+\mathrm{i}pv$ in the PL, we obtain the following relations:
$p$ $=$ $\frac{\omega}{E}-\frac{E\omega}{f_{\rm c}}\Lambda+m\beta\Lambda,$ (2.36)
$\tilde{\omega}$ $=$ $\beta\left(\frac{L}{E}\omega-m\right).$ (2.37)
@@
$\omega$ $=$ $\frac{E}{L}\left(n_{y}+m+\frac{1}{2}\right)-\mathrm{i}\frac{E}{L}\frac{\alpha}{\beta}\left(n_{x}+\frac{1}{2}\right)$ (2.38)
$=$ $\Omega\left(n_{y}+m+\frac{1}{2}\right)-\mathrm{i}\lambda\left(n_{x}+\frac{1}{2}\right).$ (2.39)
Therefore, identifying $n_{y}+m$ to $\ell$ we can establish the correspondence between the null geodesics and QNMs reported in Ref. [13]. Since $\omega$ is a complex variable, $p$ is also complex in general, and the amplitude of $\Psi$ changes along the $v$-direction unless $p$ is a real value. To make the value of $p$ real, we need to fix the arbitrary constant $\Lambda$ to be $\frac{f_{\rm c}}{E^{2}}$. Then we obtain $p=m/L$, and the relevant components of the coordinate transformation become
$t$ $=$ $\frac{E}{f_{\rm c}}u,$ (2.40)
$\phi$ $=$ $\beta u+\frac{v}{L}.$ (2.41)
@@
$P^{\ell}_{\ell}(\cos\theta)\sim\left(1-\cos^{2}\left(\frac{\pi}{2}+\frac{y}{r_{\rm c}}\right)\right)^{\ell/2},$ (2.42)
@@
$P^{\ell}_{\ell}(\cos\theta)\sim 1-\frac{\ell}{2}\frac{y^{2}}{r_{\rm c}^{2}}.$ (2.43)
A similar expression can be obtained by expanding the ground state for the SHO described by Eq. (2.30) as follows:
$\exp\left[-\frac{1}{2}\beta py^{2}\right]\sim\exp\left[-\frac{1}{2}\beta\frac{m}{L}y^{2}\right]=\exp\left[-\frac{1}{2}\frac{mE}{f_{\rm c}L}\Omega y^{2}\right]\sim 1-\frac{n_{y}+m}{2}\frac{y^{2}}{r_{\rm c}^{2}},$ (2.44)
@@
$\partial_{\theta}+\mathrm{i}\cot\theta\partial_{\phi}\sim-r_{\rm c}\partial_{y}-\cot\left(\frac{y}{r_{\rm c}}+\frac{\pi}{2}\right)m\sim-r_{\rm c}\partial_{y}+\frac{L}{r_{\rm c}}py=r_{\rm c}(-\partial_{y}+\beta py).$ (2.45)